OpenStudy (anonymous):

how do you take out the natural log? like in this case... 1=lnx i completely forgot how to do this...

OpenStudy (nikita2):

use the rule e^ln(x) = x

OpenStudy (anonymous):

so in that 1=lnx, can i just do e^1=e^lnx and then solve?

OpenStudy (nikita2):

x = e ? )

OpenStudy (anonymous):

thanks!